OpenStudy (anonymous):

H(x) = x^2 + 3 H^-1 is a function. (T/F)

hartnn (hartnn):

do you know how to get inverse of H(x) ?

OpenStudy (anonymous):

No.. not really. Do I just put it over 1?

OpenStudy (anonymous):

I mean under 1

hartnn (hartnn):

nopes, lets say H(x)=y y= x^2+3 now INTERCHANGE x and y so, x = y^2+3 now can you isolate 'y' from here ?? (that would be your inverse function)

OpenStudy (anonymous):

I would then have y^2 = x-3 and then y = sqrt(x-3)

hartnn (hartnn):

very good! just one modification, when you take the square root, add +/- sign... \(if \: a^2=b, \:\: a=\pm\sqrt b\) got this ?

hartnn (hartnn):

so, your inverse function will be \(H^{-1}(x)=\pm \sqrt{x-3}\)

hartnn (hartnn):

can you tell me whether its a function or not ?

OpenStudy (anonymous):

No because a function can't be square rooted, right?

hartnn (hartnn):

a function can be square rooted, but yes, this is not a function because 'for one value of 'x' we are getting 2 values of inverse function' (like if you put x=12, we get +3 and -3) this clearly violates definition of a function.

hartnn (hartnn):

in short, this is not a function because of that \(\huge \pm\) sign..... got this ?

OpenStudy (anonymous):

Oh, yes I do. Thank you! :) But quick question. So a function can be square rooted, but not with the +/- sign, so essentially, it can't be square rooted?

hartnn (hartnn):

well yes, you can say that. a function cannot be square rooted to get another function :)

OpenStudy (anonymous):

Okay, thank you:)

hartnn (hartnn):

welcome ^_^